# Marco Randoa1, Martin Jamesb, Alessandro Verria, Lorenzo Rosascoa, Agnese Seminarab1

### v1

@@ Results
By repeating training using different values of $T$ we find that performance depends on memory and an optimal memory $T^{*}$ exists (Figure 2A). Why is there an optimal memory? The shortest memory $T=1$ corresponds to instantaneous olfactory states: the instantaneous contour maps of the olfactory states are convoluted and the void state is pervasive (Figure 2C, top). As a consequence, agents often activate recovery even when they are within the plume, the policy almost always leads to the source ($f^{+}=79\%\pm 13\%$) but follows lengthy convoluted paths ($\tau_{\text{min}}/\tau=0.14\pm 0.05$, Figure 2A, bottom). As memory increases, the olfactory states become smoother and agents encounter less voids (Figure 2C, center), perform straighter trajectories ($\tau_{\text{min}}/\tau=0.5\pm 0.3$) and reach the source reliably ($f^{+}=95\%\pm 8\%$), Figure 2C, bottom. Further increasing memory leads to even less voids within the plume and even smoother olfactory states. However – perhaps surprisingly – performance does not further improve but slightly decreases (at $T=50$, $f^{+}=94\%\pm 8\%$ and $\tau_{\text{min}}/\tau=0.38\pm 0.36$). A long memory is deleterious because it delays recovery from accidentally exiting the plume, thus increases the number of voids outside of the plume (Figure 2C, bottom). Indeed, agents often leave the plume accidentally as they measure their olfactory state while they move. They receive no warning, but realize their mistake after $T$ steps, when they enter the void state and activate recovery to re-enter the plume. The delay is linear with memory when agents recover by backtracking, but it depends on the recovery strategy (see discussion below and Supplementary Figure 1).
@@
There is no way to select the optimal memory $T^{*}$ without comparing several agents or relying on prior information on the blank durations. In order to avoid prior information, we venture to define memory adaptively along each agent’s path, using the intuition outlined above. We define a buffer memory $T_{b}$, and let the agent respond to a sensing window $T<T_{b}$. Ideally we would like to set $T\sim\langle\tau_{b}\rangle$. With this choice, blanks shorter than the average blank are ignored, as they are expected within the plume, whereas blanks longer than average initiate recovery, as they signal that the agent exited the plume. However, agents do not have access to $\langle\tau_{b}\rangle$ hence we set $T=\tau_{b}^{-}$, where $\tau_{b}^{-}$ is the most recent blank experienced by the agent. With this choice, the sensing memory $T$ fluctuates considerably along an agent’s path, due to turbulence ([32] and Figure 3A-B). Note that blanks are estimated along paths, thus the statistics of $T$ only qualitatively matches the Eulerian statistics of $\tau_{b}$. Despite the fluctuations, performance using the adaptive memory nears performance with the optimal memory (Figure 3C). This result confirms our intuition that memory should match the blank time. The advantage of the adaptive memory is that it relies solely on experience, with no prior information whatsoever. This is unlike $T^{*}$ which can only be selected using prior information, with no guarantee of generalization to other plumes.
@@
Learning to recover.
So far, our agents combine a learned policy from non-void states to a heuristics from the void state, which we called the recovery strategy. We have considered a biologically-inspired heuristics where searchers make it back to locations within the plume by retracing their path backward. Similar results are obtained for Brownian recovery with a different optimal memory (see Materials and Methods and Supplementary Figure 1). To further strip the algorithm of heuristics, we ask whether the recovery strategy may be learned, rather than fixed a priori. To this end, we split the void state in many states, labeled with the time elapsed since first entering the void. The counter is reset to 00 whenever the searcher detects the odor. The definition of the 15 non-void states $o_{1},...,o_{15}$ remains unaltered. Interestingly, with this added degree of freedom, the agent learns an even better recovery strategy as reflected by all our measures of performance Figure 3D. Note that the learned recovery strategy resembles the casting behavior observed in flying insects [37], as discussed below.
Figure 5: Generalization to statistically different environments. (A) Snapshots of odor concentration normalizes with concentration at the source, colorcoded from blue (0) to yellow (1) for environment 1 to 6 from top to bottom. Environment $1^{*}$ is the native environment where all agents are trained. (B) Performance for the three recovery strategies backtracking (green) Brownian (red) and learned (blue), with adaptive memory, trained on the native environment and tested across all environments. Four measures of performance defined in the main text are shown. Dark squares mark the mean, and empty rectangles $\pm$ standard deviation. No standard deviation is shown for the $f^{+}$ measure in the learned case as this strategy is deterministic (see Materials and Methods).
@@
Characterization of the optimal policies. To understand how different recoveries affect the agent’s behavior, we characterize the optimal policies obtained using the three recovery strategies. We visualize the probability to encounter each of the 16 olfactory states, or occupancy (circles in Figure 4), and the spatial distribution of the olfactory states. In the void state, the agent activates the recovery strategy. Recovery from the void state affects non-void olfactory states as well: their occupancy, their spatial distribution, and the action they elicit (Figure 4 and Supplementary Figure 3). This is because the agent computes its olfactory state online, according to its prior history which is affected by encounters with the void state. However, for all recoveries, non-void states are mostly encountered within the plume and largely elicit upwind motion (Figure 4, top, center). Thus macroscopically, all agents learn to surge upwind when they detect any odor within their memory, and to recover when their memory is empty. This suggests a considerable level of redundancy which may be leveraged to reduce the number of olfactory states, thus the computational cost.
@@
The void state shows the most relevant differences: for both heuristic recoveries, 40% or more of the agents are in the void state and they are spatially spread out. In contrast, in the case of learned recovery, the optimal policy limits occurrence of the void state to 26% of the agents, confined to a narrow band near the edge of the plume. From these locations, the agents quickly recover the plume, explaining the boost in performance discussed above. Note that, exclusively for the learned recovery, the optimal policy is enriched in actions downwind to avoid overshooting the source. Indeed, from positions beyond the source, the learned strategy is unable to recover the plume as it mostly casts sideways, with little to no downwind action. Intuitively, the precise locations where agents move downwind may be crucial to efficiently avoid overshooting. Thus the policy may depend on specific details of the odor plume, consistent with poorer generalization of the learned recovery (discussed next).
@@
Finally, we test performance of the trained agents on six environments, characterized by distinct fluid flows and odor plumes (Figure 5 and Materials and Methods). Environment 1 is the native environment, where the agents were originally trained; Environment 2 is obtained by increasing the threshold of detection, which makes the signals considerably more sparse with longer blanks. Environments 3 and 4 are closer to the lower surface of the simulated domain, where the plume is smaller and fluctuates less. Environment 5 is a similar geometry, but obtained for a smaller Reynolds number and a different way to generate turbulence. Finally Environment 6 has an even larger Reynolds number, a longer domain and a smaller source, which creates an even more dramatically sparse signal. We consider agents with adaptive memory and compare the three recovery strategies discussed above – Brownian, backtracking and learned, see Figure 5B. Comparing performance across environments we find that: (i) although performance is degraded when testing in non native environments, all agents with adaptive memory are still extremely likely to find the source; (ii) Brownian recovery has lowest performance and generalization across all environments; (iii) backtracking provides good performance and generalization; (iii) the learned recovery strategy performs best in all environments by all performance metrics. In the most intermittent Environment 6 a striking 91% of agents succeeds in finding the source, with trajectories less the twice as long as the shortest path to the source. The upshot of generalization is that agents may navigate distinct turbulent plumes using a baseline strategy learned in a specific plume. Importantly, even if performance (mildly) degrades, most agents still do reach the source, suggesting that fine-tuning this strategy may enable efficient adaptation to different environments. Further work is needed to establish how much fine-tuning is needed to fully adapt the baseline strategy to different environments.
Discussion
@@
In this work, we showed that agents exposed to a turbulent plume learn to associate salient features of the odor time trace – the olfactory state – to an optimal move that guides them to the odor source. The upshot of responding solely to odor is that the agent does not navigate based on where it believes the target is and thus needs no map of space nor prior information about the odor plume, which avoids considerable computational burden. On the flip side, in our stimulus-response algorithm, agents need to start from within the plume, however sparse and fragmented. Indeed, far enough from the source, Q-learning agents are mostly in the void state and they can only recover the plume if they have previously detected the odor or are right outside the plume. In contrast, agents using a map of space can navigate from larger distances than are reachable by responding directly to odor cues. Indeed, in the map-based POMDP setting, absence of odor detection is still informative and it enables agents to first find the plume, and then refine the search to localize the target within the plume [26, 38].
@@
We show that because turbulent odor plumes constantly switch on and off, navigation must handle both absence and presence of odor stimuli. We address this fundamental issue by alternating between two distinct strategies: (i) Prolonged absence of odor prompts entry in the void state and triggers a recovery strategy to make contact with the plume again. We explored two heuristic recoveries and found that back-tracking to where the agent last detected odor is much more efficient than Brownian recovery. An even more efficient recovery can be learned that resembles cross-wind casting and limits the void state to a narrow region right outside of the plume. Casting is a well-studied computational strategy [12, 5] also observed in animal behavior, most famously in flying insects [37]. Intriguingly, cast and surge also emerges in algorithms making use of a model of the odor, whether for Bayesian updates or for policy optimization [3, 26, 30]. Whether natural casting behavior is learned, as in Q-learning, or is hard-wired in a model of the odor plume remains a fascinating question for further research. (ii) Odor detections prompt entry in non-void olfactory states, which predominantly elicit upwind surge. Blanks shorter than the sensing memory are ignored, i.e. agents do not enact recovery but respond to stimuli experienced prior to the short blank. Further work may optimize these non-void olfactory states by feature engineering, e.g. testing different discretizations to reduce redundancy or screening a large library of features using supervised learning as in [33] to potentially improve performance. Alternatively, feature engineering may be bypassed altogether by the use of recurrent neural networks (RNNs) as recently proposed in [29], possibly at the expense of interpretability. A systematic comparison using a common dataset is needed to elucidate how other heuristic and normative model-free algorithms handle odor presence vs odor absence.
@@
Thus in Q-learning, memory is a temporal window matching odor blanks and distinguishing whether agents are in or out of the plume. The role of memory for olfactory search has been recently discussed in ref. [30]. In POMDPs, memory is stored in a detailed belief of agent position relative to the source. In finite state controllers, memory denotes an internal state of the agent and was linked to a coarse grained belief of the searcher being within or outside of the plume, similar to our findings. In recurrent neural networks memory is stored in the learned weights. A quantitative relationship between these different forms of memory and their connection to spatial perception remains to be understood.
@@
We conclude by listing a series of experiments to test these ideas in living systems. First, olfactory search in living systems displays memory ([30, 10] and refs. therein). In insects, temporal scales can be measured associated to memory. Indeed, for flying insects loss of contact with a pheromone plume triggers crosswind casting and sometimes even downwind displacement [7, 39]. Interestingly, the onset of casting is delayed with respect to loss of contact with the plume, but this delay is not understood [39, 40]. In walking flies, the timing of previous odor encounters biases navigation [41]. (How) do these temporal timescales depend on the waiting times between previous detections? Using optogenetics [42, 43, 44, 45] or olfactory virtual reality with controlled odor delivery [46] experiments may measure memory as a function of the full history of odor traces. For insects, one may monitor memory by tracking the onset of cross wind casting with respect to the loss of the plume. More in general, a temporal memory may be defined by monitoring how far back in the past should two odor traces be identical in order to elicit the same repertoire of motor controls.
@@
Second, our algorithm learns a stimulus-response strategy that relies solely on odor cues. The price to pay is that the agent must follow the ups and downs of the odor trace in order to compute averages and recognize blanks. A systematic study may use our algorithm to test the requirements of fidelity of this temporal representation, and how it depends on turbulence. How does turbulence affect the fidelity of odor temporal representation in living systems? Crustaceans provide excellent model system to ask this question, as they are known to use bursting olfactory receptor neurons to encode temporal information from olfactory scenes [47, 48]. Temporal information is also encoded in the olfactory bulb of mammals [49, 50]. Organisms with chemo-tactile systems like the octopus [51] may serve as a comparative model, to ask whether touch-chemosensation displays a sloppier temporal response, reflecting that surface-bound stimuli are not intermittent.
@@
Third, our Q-learning algorithm requires the agent to receive olfactory information, thus start near or within the odor plume. In contrast, algorithms making use of a spatial map and prior information on the odor plume may first search for the plume (in conditions of near zero information) and then search the target within the plume [26, 38, 3]. Animals are known to use prior information to home into regions of space where the target is more likely to be found; but they can switch to navigation in response to odor (see e.g. [7, 8, 9, 10]). What triggers the switch from navigation driven by spatial perception to navigation driven by odor? For mice, the need for spatial perception may be tested indirectly by comparing paths in light vs dark, noting that neuronal place fields, that mediate spatial perception, are better stabilized by vision than olfaction [52, 53]. Thus in the light, animals have the ability to implement both map-less and map-based algorithms, whereas in the dark they are expected to more heavily rely on map-less algorithms. To make sure animals start searching for the odor target even before sensing odor, operant conditioning can be deployed so that animals associate an external cue (e.g. a sound) to the beginning of the task.
@@
Acknowledgements
This research was supported by grants to A.S. from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002724 RIDING), the Air Force Office of Scientific Research under award number FA8655-20-1-7028, and the National Institutes of Health (NIH) under award number R01DC018789. L.R. acknowledges the financial support of the European Research Council (grant SLING 819789). We thank Francesco Viola for support and discussions regarding computational fluid dynamics as well as Antonio Celani, Venkatesh Murthy, Yujia Qi, Francesco Boccardo, Luca Gagliardi, Francesco Marcolli and Arnaud Ruymaekers for comments on the manuscript.
Materials and Methods

### v2

@@ 2 Results
[…]
[Optimal memory with Brownian recovery]The role of temporal memory with Brownian recovery strategy (same as main Figure 2A). (A) Total cumulative reward (top left) and standard deviation (top right) as a function of memory showing an optimal memory $T^{*}=3$ for the Brownian agent. Other measures of performance with their standard deviations show the same optimal memory (bottom). The trade-off between long and short memories discussed in the main text holds, but here exiting the plume is much more detrimental because regaining position within the plume by Brownian motion is much lengthier. (B) As for the Backtracking agent, the optimal memory corresponds minimizes the number of times the agent encounters the void state.Refer to caption
By repeating training using different values of $T$ we find that performance depends on memory and an optimal memory $T^{*}$ exists (Figure 2A). Why is there an optimal memory? The shortest memory $T=1$ corresponds to instantaneous olfactory states: the instantaneous contour maps of the olfactory states are convoluted and the void state is pervasive (Figure 2C, top). As a consequence, agents often activate recovery even when they are within the plume, the policy almost always leads to the source ($f^{+}=79\%\pm 13\%$) but follows lengthy convoluted paths ($\tau_{\text{min}}/\tau=0.14\pm 0.05$, Figure 2C, bottom). As memory increases, the olfactory states become smoother and agents encounter less voids (Figure 2C, center), perform straighter trajectories ($\tau_{\text{min}}/\tau=0.5\pm 0.3$) and reach the source reliably ($f^{+}=95\%\pm 8\%$), Figure 2A, bottom. Further increasing memory leads to even less voids within the plume and even smoother olfactory states. However – perhaps surprisingly – performance does not further improve but slightly decreases (at $T=50$, $f^{+}=94\%\pm 8\%$ and $\tau_{\text{min}}/\tau=0.38\pm 0.36$). A long memory is deleterious because it delays recovery from accidentally exiting the plume, thus increases the number of voids outside of the plume (Figure 2C, bottom). Indeed, agents often leave the plume accidentally as they measure their olfactory state while they move. They receive no warning, but realize their mistake after $T$ steps, when they enter the void state and activate recovery to re-enter the plume. The delay is linear with memory when agents recover by backtracking, but it depends on the recovery strategy (see Materials and Methods and Figure 2–figure supplement 1).
@@
There is no way to select the optimal memory $T^{*}$ without comparing several agents or relying on prior information on the blank durations. In order to avoid prior information, we venture to define memory adaptively along each agent’s path, using the intuition outlined above. We define a buffer memory $T_{b}$, and let the agent respond to a sensing window $T<T_{b}$. Ideally we would like to set $T\sim\langle\tau_{b}\rangle$. With this choice, blanks shorter than the average blank are ignored, as they are expected within the plume, whereas blanks longer than average initiate recovery, as they signal that the agent exited the plume. However, agents do not have access to $\langle\tau_{b}\rangle$ hence we set $T=\tau_{b}^{-}$, where $\tau_{b}^{-}$ is the most recent blank experienced by the agent. With this choice, the sensing memory $T$ fluctuates considerably along an agent’s path, due to turbulence (Celani et al. (2014) and Figure 3A-B). Note that blanks are estimated along paths, thus the statistics of $T$ only qualitatively matches the Eulerian statistics of $\tau_{b}$. Despite the fluctuations, performance using the adaptive memory nears performance with the optimal memory (Figure 3C). This result confirms our intuition that memory should match the blank time. The advantage of the adaptive memory is that it relies solely on experience, with no prior information whatsoever. This is unlike $T^{*}$ which can only be selected using prior information, with no guarantee of generalization to other plumes.
Figure 3: The adaptive memory approximates the duration of the blank dictated by physics and it is an efficient heuristics, especially when coupled with a learned recovery strategy. (A) Top to bottom: Colormaps of the Eulerian average blank time $\tau_{b}$; average sensing memory $T$; standard deviation of Eulerian blank time and of sensing memory. The sensing memory statistics is computed over all agents that are located at each discrete cell, at any point in time. (B) Probability distribution of $\tau_{b}$ across all spatial locations and times (black) and of $T$ across all agents at all times (gray). (C) Performance with the adaptive memory nears performance of the optimal fixed memory, here shown for backtracking; similar results apply to the Brownian recovery (Figure 3–figure supplement 2). (D) Comparison of five recovery strategies with adaptive memory: The learned recovery with adaptive memory outperforms all fixed and adaptive memory agents. In (C) and (D) dark squares mark the mean, and light rectangles mark $\pm$ standard deviation. $f^{+}$ is defined as the fraction of agents that reach the target at test, hence has no standard deviation.
[…]
[All performance metrics with fixed and adaptive memory and two different recovery strategies.]All four measures of performance across agents with fixed memory and Backtracking vs Brownian recovery (green and red respectively, unframed boxes) and with adaptive memory for Backtracking, Brownian and Learned recovery (green, red and blue respectively, framed boxes)Refer to caption
Learning to recover.
So far, our agents combine a learned policy from non-void states to a heuristics from the void state, which we called the recovery strategy. We have considered a biologically-inspired heuristics where searchers make it back to locations within the plume by retracing their path backward. To further strip the algorithm of heuristics, we ask whether the recovery strategy may be learned, rather than fixed a priori. To this end, we split the void state in many states, labeled with the time elapsed since first entering the void. We pick 50 void states, as less than 50 void states results in no convergence, and states above 50 are useless because rarely visited. The counter is reset to 00 whenever the searcher detects the odor. The definition of the 15 non-void states $o_{1},...,o_{15}$ remains unaltered. Interestingly, with this added degree of freedom, the agent learns an even better recovery strategy as reflected by all our measures of performance Figure 3D. Note that the learned recovery strategy resembles the casting behavior observed in flying insects David et al. (1983), as discussed below. In fact, insects deploy a range of recovery strategies depending on locomotor mode and environment. To corroborate these results, we compare performance using two additional biologically-inspired recovery strategies, i.e. circling (observed in windless environments Stupski and van Breugel (2024)), and cast & surge David et al. (1983) as well as a Brownian recovery which does not have a direct biological relevance but represents a simple computational benchmark. The learned recovery outperforms all heuristic recoveries, as seen by the cumulative reward $G$ (Figure 3D). Circling is the second best recovery and shortly follows the learned recovery. Circling achieves nearly optimal performance by further decreasing failures (metrics $f^{+}$), but slowing down (metrics $g^{+}$ and $\tau_{min}/\tau$).
Characterization of the optimal policies.
@@
[…]
@@
To understand how different recoveries affect the agent’s behavior, we characterize the optimal policies obtained using the three recovery strategies. We visualize the probability to encounter each of the 16 olfactory states, or occupancy (circles in Figure 4), and the spatial distribution of the olfactory states. In the void state, the agent activates the recovery strategy. Recovery from the void state affects non-void olfactory states as well: their occupancy, their spatial distribution, and the action they elicit (Figure 4 and Figure 4–figure supplement 3). This is because the agent computes its olfactory state online, according to its prior history which is affected by encounters with the void state. However, for all recoveries, non-void states are mostly encountered within the plume and largely elicit upwind motion (Figure 4, top, center). Thus macroscopically, all agents learn to surge upwind when they detect any odor within their memory, and to recover when their memory is empty. This suggests a considerable level of redundancy which may be leveraged to reduce the number of olfactory states, thus the computational cost. Reducing the number of non-empty olfactory states drastically to just 1 does indeed show degraded performance (see Figure 5 - figure supplement 5). A systematic optimization of odor representation requires a considerable reformulation of the algorithm which is beyond the scope of the current work. Note that, exclusively for the learned recovery, the optimal policy is enriched in actions downwind to avoid overshooting the source. Indeed, from positions beyond the source, the learned strategy is unable to recover the plume as it mostly casts sideways, with little to no downwind action. Intuitively, the precise locations where agents move downwind may be crucial to efficiently avoid overshooting. Thus the policy may depend on specific details of the odor plume, consistent with poorer generalization of the learned recovery (discussed next). We expect that in conditions where overshooting the source is more prominent, downwind motion may emerge as an effective component of the recovery strategy, similar to observations in insects (e.g. Wolf and Wehner (2000); Alvarez-Salvado et al. (2018).
@@
The void state shows the most relevant differences: for both heuristic recoveries, 40% or more of the agents are in the void state and they are spatially spread out. In contrast, in the case of learned recovery, the optimal policy limits occurrence of the void state to 26% of the agents, confined to a narrow band near the edge of the plume. From these locations, the agents quickly recover the plume, explaining the boost in performance discussed above. In all distinct trainings of the agents with learned-recovery, we observed that the trajectories in the void start with an initial surge of 6$\pm$2 steps; continue with either crosswind direction and then switch to the other side, with little to no downwind actions (see Figure 4 - Figure Supplement 4 and Table 1). This recovery thus mixes aspects of exploitation (surge) to aspects of exploration (cast): we defer a more in-depth analysis that disentangles these two aspects elsewhere.
Table 1: Parameters of the learned recovery, statistics over 20 independent trainings.
Initial surge upwind $6\pm 2$
Total steps upwind $15\pm 2$
Total steps downwind $1.3\pm 1.4$
Total steps to the right $15\pm 3$
Total steps to the left $18\pm 6$
@@
Finally, we test performance of the trained agents on six environments, characterized by distinct fluid flows and odor plumes (Figure 5 and Materials and Methods). Environment 1 is the native environment, where the agents were originally trained; Environment 2 is obtained by increasing the threshold of detection, which makes the signals considerably more sparse with longer blanks. Environments 3 and 4 are closer to the lower surface of the simulated domain, where the plume is smaller and fluctuates less. Environment 5 is a similar geometry, but obtained for a smaller Reynolds number and a different way to generate turbulence. Finally Environment 6 has an even larger Reynolds number, a longer domain and a smaller source, which creates an even more dramatically sparse signal. Environments 1, 2, 5 and 6 are representative of conditions encountered far from a substrate, e.g. by flying or swimming organisms. Environments 3 and 4 represent odor near surfaces relevant to terrestrial or benthic navigation (but not directly applicable to trail tracking, where odor traces on the substrate is tracked). Note that we consider a Schmidt number Sc = 1 appropriate for odors in air but not in water. However, we expect a weak dependence on the Schmidt number as the Batchelor and Kolmogorov scales are below the size of the source and we are interested in the large scale statistics Falkovich et al. (2001); Celani et al. (2014); Duplat et al. (2010).
@@
We consider agents with adaptive memory and compare the five recovery strategies discussed above – backtracking, learned, circling, cast & surge and Brownian, see Figure 5B. Comparing performance across environments we find that: (i) although performance is degraded when testing in non native environments, backtracking, learned and circling recoveries with adaptive memory are still extremely likely to find the source. The upshot of generalization is that agents may navigate distinct turbulent plumes using a baseline strategy learned in a specific plume. Importantly, as most of these agents still do reach the source, fine-tuning may enable efficient adaptation to different environments. Further work is needed to establish how much fine-tuning is needed to fully adapt to different environments. (ii) Brownian and cast& surge recoveries have lowest performance and generalization across all environments. Cast & surge is often used as a comparison Verano et al. (2023) and can be extremely effective: our results do not contradict the literature, but simply showcase that cast width and surge length need to be carefully defined. (iii) The cumulative reward $G$ shows that the learned recovery is the best at generalizing to environments 2, 5 and 6. Particularly, in the most intermittent Environment 6 a striking 91% of agents succeeds in finding the source, with trajectories less the twice as long as the shortest path to the source. (iv) Circling is the best at generalizing to environments 3 and 4, representative of less intermittent regions near the substrate. As observed for the native environment, circling favors success rate (metrics $f^{+}$) against speed (metrics $g^{+}$ and $\tau_{\text{min}}/\tau$).
Figure 5: Generalization to statistically different environments. (A) Snapshots of odor concentration normalized with concentration at the source, colorcoded from blue (0) to yellow (1) for environment 1 to 6 as labeled. Environment $1^{*}$ is the native environment where all agents are trained. (B) Performance for the five recovery strategies brownian (red), backtracking (green), learned (blue), circling (orange) and zigzag (purple) with adaptive memory, trained on the native environment and tested across all environments 1 to 6. Four measures of performance defined in the main text are shown. Dark squares mark the mean, and empty rectangles $\pm$ standard deviation. No standard deviation is shown for the $f^{+}$ measure for the learned, circling and zigzag recoveries as these strategies are deterministic (see Materials and Methods).
[…]
3 Discussion
@@
In this work, we showed that agents exposed to a turbulent plume learn to associate salient features of the odor time trace – the olfactory state – to an optimal move that guides them to the odor source. The upshot of responding solely to odor is that the agent does not navigate based on where it believes the target is and thus needs no map of space nor prior information about the odor plume, which avoids considerable computational burden. The only spatial awareness needed to implement this algorithm is the ability to orient motion relative to the mean flow which is assumed known. In reality animals cannot measure the mean flow but rely on local measures of flow speed, using e.g. antennas for insects Reynolds et al. (2010), whiskers for rodents Yu et al. (2016) or the lateral line for marine organisms Liao (2006). Further work is needed to bridge the gap with our simplified setting. On the flip side, in our stimulus-response algorithm, agents need to start from within the plume, however sparse and fragmented. Indeed, far enough from the source, Q-learning agents are mostly in the void state and they can only recover the plume if they have previously detected the odor or are right outside the plume. In contrast, agents using a map of space can navigate from larger distances than are reachable by responding directly to odor cues. Indeed, in the map-based POMDP setting, absence of odor detection is still informative and it enables agents to first find the plume, and then refine the search to localize the target within the plume Rigolli et al. (2022b); Loisy and Heinonen (2023a).
@@
We show that because the odor signal within a turbulent plume constantly switches on and off, navigation must handle both absence and presence of odor stimuli. We address this fundamental issue by alternating between two distinct strategies: (i) Prolonged absence of odor prompts entry in the void state and triggers a recovery strategy to make contact with the plume again. We explored four heuristic recoveries and found that spiralling around the location where the agent last detected odor is the most efficient heuristics, which privileges reliable success rather than speed. An even more efficient recovery can be learned that resembles cross-wind casting and limits the void state to a narrow region right outside of the plume. Casting is a well-studied computational strategy Baker (1990); Balkovsky and I. (2002) also observed in animal behavior, most famously in flying insects David et al. (1983). Intriguingly, cast and surge also emerges in algorithms making use of a model of the odor, whether for Bayesian updates or for policy optimization Vergassola et al. (2007); Rigolli et al. (2022b); Verano et al. (2023). Whether natural casting behavior is learned, as in Q-learning, or is hard-wired in a model of the odor plume remains a fascinating question for further research. Clearly, the width of the casts and length of the surges are of crucial importance: a hard-wired cast and surge recovery with arbitrary parameters shows poor performance. (ii) Odor detections prompt entry in non-void olfactory states, which predominantly elicit upwind surge. Blanks shorter than the sensing memory are ignored, i.e. agents do not enact recovery but respond to stimuli experienced prior to the short blank. The non-void olfactory states are crafted based on biologically-plausible features which have been shown to harness positional information. Further work may optimize these non-void olfactory states by feature engineering, e.g. testing different discretizations to reduce redundancy. A drastic reduction to a single non-void olfactory state degrades performance, suggesting screening a large library of features using supervised learning as in Rigolli et al. (2022a) may be used to potentially improve performance. This approach will provide a systematic ranking of the most efficient temporal features of odor time traces for navigation; however it will have to test different memories, discretizations and regression algorithms as well, making it cumbersome. Alternatively, feature engineering may be bypassed altogether by the use of recurrent neural networks (RNNs) Singh et al. (2023) or finite state controllers Verano et al. (2023). These algorithms are appealing in that they bypass entirely the need to hand-craft explicit features of odor time traces. On the flip side, they provide no explicit handle at the odor features that drive behavior nor at the specific duration of the temporal memory and how it is related to the physics of the odor cues. Thus to extract these information extra work is needed to interrogate these algorithms. For example, principal component analysis in Singh et al. (2023) suggests the hidden state of trained agents correlates with biologically relevant variables, including head direction, odor concentration and time since last detection. Finally, a systematic comparison using a common dataset is needed to elucidate how other heuristic and normative model-free algorithms handle odor presence vs odor absence.
@@
Thus in Q-learning, memory is a temporal window matching odor blanks and distinguishing whether agents are in or out of the plume. The role of memory for olfactory search has been recently discussed in ref. Verano et al. (2023). In POMDPs, memory is stored in a detailed belief of agent position relative to the source. In finite state controllers, memory denotes an internal state of the agent and was linked to a coarse grained belief of the searcher being within or outside of the plume, similar to our findings. In recurrent neural networks memory is stored in the learned weights. A quantitative relationship between these different forms of memory and their connection to spatial perception remains to be understood.
@@
We conclude by listing a series of experiments to test these ideas in living systems. First, olfactory search in living systems displays memory (Verano et al. (2023); Baker et al. (2018) and refs. therein). In insects, temporal scales can be measured associated to memory. Indeed, for flying insects loss of contact with a pheromone plume triggers crosswind casting and sometimes even downwind displacement Cardé (2021); Kuenen and Cardé (1994). Interestingly, the onset of casting is delayed with respect to loss of contact with the plume Kuenen and Cardé (1994); van Breugel F and MH (2014), but this delay is not understood. Similarly, an odor detection elicits upwind surges that can last several seconds Kathman et al. (2024); Alvarez-Salvado et al. (2018). In walking flies, the timing of previous odor encounters biases navigation Demir et al. (2020b). (How) do these temporal timescales depend on the waiting times between previous detections? Using optogenetics Gepner et al. (2015); Hernandez-Nunez et al. (2015); Matheson et al. (2022); David Stupski and van Breugel (2024) or olfactory virtual reality with controlled odor delivery Radvansky and Dombeck (2018) experiments may measure memory as a function of the full history of odor traces. For insects, one may monitor memory by tracking the onset of cross wind casting with respect to the loss of the plume. More in general, a temporal memory may be defined by monitoring how far back in the past should two odor traces be identical in order to elicit the same repertoire of motor controls.
@@
Second, our algorithm learns a stimulus-response strategy that relies solely on odor cues. The price to pay is that the agent must follow the ups and downs of the odor trace in order to compute averages and recognize blanks. A systematic study may use our algorithm to test the requirements of fidelity of this temporal representation, and how it depends on turbulence. How does turbulence affect the fidelity of odor temporal representation in living systems? Crustaceans provide excellent model system to ask this question, as they are known to use bursting olfactory receptor neurons to encode temporal information from olfactory scenes Bobkov and Ache (2007); Ache et al. (2016). Temporal information is also encoded in the olfactory bulb of mammals Carey et al. (2009); Ackels et al. (2021a). Organisms with chemo-tactile systems like the octopus Allard et al. (2023) may serve as a comparative model, to ask whether touch-chemosensation displays a sloppier temporal response, reflecting that surface-bound stimuli are not intermittent.
@@
Third, our Q-learning algorithm requires the agent to receive olfactory information, thus start near or within the odor plume. In contrast, algorithms making use of a spatial map and prior information on the odor plume may first search for the plume (in conditions of near zero information) and then search the target within the plume Rigolli et al. (2022b); Loisy and Heinonen (2023a); Vergassola et al. (2007). Animals are known to use prior information to home into regions of space where the target is more likely to be found; but they can switch to navigation in response to odor (see e.g. Cardé (2021); Schal (1982); Gire et al. (2016); Baker et al. (2018)). What triggers the switch from navigation driven by spatial perception to navigation driven by odor? For mice, the need for spatial perception may be tested indirectly by comparing paths in light vs dark, noting that neuronal place fields, that mediate spatial perception, are better stabilized by vision than olfaction Save et al. (2000); Zhang and D (2015). Thus in the light, animals have the ability to implement both map-less and map-based algorithms, whereas in the dark they are expected to more heavily rely on map-less algorithms. To make sure animals start searching for the odor target even before sensing odor, operant conditioning can be deployed so that animals associate an external cue (e.g. a sound) to the beginning of the task. Note that distinct species control locomotion differently and as a result trajectories are usually far more complex than a sequence of discrete steps on a checker board. Thus to compare algorithms to animal behavior a more detailed model of the specific motor controls is to be developed.
@@
4 Methods and Materials